INAF queue observing, 2018 Feb 03-04

Observer: F. Cusano, A. Gargiulo
Telescope Operator: D. Gonzalez-Huerta
Support Astronomer: B. Rothberg
Instrument:

Summary

At the sunset the sky is clear
instruments: LBC works, LUCI1-2 are ok, MODS2 blue grating only
problem with MOS exchange of LUCI1 at 4:28, we can not use LUCI1 for the rest of the night and we can not do AO

Night duration (12\xB0 el twilight) 2:00am- 13:15pm UT = 11 h 15m

Observing time: 10h 55m (97%)

Weather time loss: 0h m (0%)*

technical time loss h 20*m* (3%)

Data Summary

Project_RUN ObjectsSorted descending Instrument Filter ExpTime Notes
Prandoni USS415 LUCI2 K,J 0.6h completed, in J guide probe in the fov
Prandoni USS36 LUCI2 J 0.2h completed
LBC flat sky flat LBC B,V    
Annibali LBC R3 LBC g,r 0.33h completed
Cusano NW-1,2,3 LBC B, V 1h 3m up to sequence 17
Gilli J LUCI2 200g HKspec 3h  
LUCI2 cal flat LUCI2 J,K   imaging
Saracco C1_23152 LUCI2 200g HKspec 2h  
LBC cal bias LBC      
LUCI2 cal arc, flat Saracco LUCI2 200g HKspec   slit 0.75"
LUCI2 cal arc, flat Gilli LUCI2 200g HKspec   slit 1"

Details

1:21 starting sky flat with LBC
  • skyflat B, V

1:34 slewing to copointing ob Cusano, NW-1
  • 1:40 running dohybrid
  • error running again at 1:43, dohybrid finished at 1:51
  • 1:54 lbcrangebal

1:56 slewing to Cusano NW-1 science
  • exec NW-1_15.xml, at 2:05 NW-2_15, at 2:14 NW-3_15
  • seeing 1.3" on images at 2:06
  • seeing 1.4" on images at 2:14
  • 2:23 exec NW-1_16
  • 2:32 dofpia
  • exec NW-2_16 at 2:34, dimm seeing 1.2" at 2:38, exec NW-3_16 at 2:44, seeing 1.3" on images at 2:46, dimm seeing 1.1" at 2:49
  • exec NW-1_17 at 2:52, dimm 1" at 2:58, seeing 1.2" on images, exec NW-2_17 at 3:01, exec NW-3_17 at 3:11
  • dimm seeing 1.2" at 3:18

3:21 slewing to Annibali R3
  • dofpia 3:22
  • 3:26 exec UGC1281_3.ob
  • dimm seeing 1.1" at 3:27, seeing 1" on images at 3:38

3:51 switch to LUCI1-2, seeing is stable at 1"

4:24 slewing to Vanzella AO JDz11
  • dimm seeing 0.9" at 4:24
  • 4:28 mos error, the SA is checking, 4:34 for the moment we can not use LUCi1, we go to LUCI2

4:37 slewing to Prandoni 36 J, LUCI2
  • 4:38 preset failed the TO is checking the pointing
  • 4:47 send again the script
  • 4:49 seeing 1.1" on guider
  • first image luci2.20180204.0001.fits
  • 5:02 seeing 0.9" on guider

5:06 communication from D. Thompson LUCI1 will not be available for the rest of the night, we keep going with LUCI2

5:08 slewing to telluric of Saracco
  • acq luci2.20180204.0013-15
  • spec 16-19

5:20 slewing to science of Saracco Target_C1_23152
  • seeing 0.7" on guider at 5:26
  • acq 20-23
  • spec 24-47
  • 5:36 seeing 0.9" on guider
  • 5:57 seeing 0.8" on guider
  • 6:32 seeing 0.7" on guider
  • 6:43 sending again the spec script
  • 7:06 seeing 1.1" on guider
  • 7:11 seeing 0.8" on guider

07:53 slewing to telluric of Gilli J, Hip 56736
  • acq 48-50
  • 8:02 seeing 0.8" on guider
  • spec 51-54

8:06 slewing to Gilli J, science
  • acq 55-58
  • 8:21 seeing 0.7" on guider
  • spec 59-112
  • 8:42 seeing 0.8" on guider
  • 9:03 seeing 0.6" on guider
  • 9:34 sending again the spec script, seeing 0.6" on guider
  • 9:51 the seeing is getting worse now is 1.1" on guider
  • 9:58 seeing 1" on guider
  • 10:04 seeing back to 0.7"
  • 10:17 seeing 0.8" on guider
  • 10:38 seeing 1.0" on guider
  • 10:46 sending again the spec script, seeing 0.9" on guider
  • 11:07 seeing 0.9" on guider
  • 11:39 seeing 0.9" on guider

11:57 slewing to telluri for Gilli J, HIP 56736
  • acq 113-115
  • spec 116-119

12:10 slewing to Prandoni USS415 K
  • 12:13 seeing 0.8"
  • first image luci2.20180204.0120.fits
  • 12:20 seeing 0.6"
  • 12:46 seeing 1.0"

12:56 sending J filter of USS415
  • 12:57 seeing 0.9"
  • guider probe in the field

* starting LUCI2 calibration

* Prandoni imaging K, J
  • 157-166 J band on/off sequence, guide inside the field we send again
  • 167-176 J band on/of sequence
  • 177-186 K band on/of sequence

*LBC bino bias

* Gilli calibration flat, arc (slit 1", 200g HKspec)
  • 187-202 on/off sequence

* Saracco calibrations flat, arc (slit 0.75", 200g HKspec)
  • 203-218 on/off sequence
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback